OpenStudy (anonymous):

solve for x when x is a real number 2x^2+5x=3

OpenStudy (anonymous):

Subtract 3 over to the otherside so we get \[2x ^{2} + 5x - 3 = 0\] From here do the quadratic formula using the given. This will solve for x.

OpenStudy (anonymous):

Using the equation you found ^

OpenStudy (acxbox22):

factor the expression above and you get: (2x -1 )(x +3 )=0 now, 2x-1=0 and x+3=0 so... x=1/2 and -3

OpenStudy (anonymous):

-b±√b^2-4ac/2a I'm solving it now hold on.

OpenStudy (acxbox22):

you dont have to use quadratic formula if you can factor and solve

OpenStudy (anonymous):

Both of our methods work, Acxbox's method is better when you can factor. The quadratic equation is better when you can't factor it.

OpenStudy (anonymous):

I just don't see how to factor it out that way.

OpenStudy (anonymous):

wait... dont you multiply the 2 to the 3 in the back and make it 6. so originally it would be 2x^2+5x-6 and you would factor it then?

OpenStudy (anonymous):

No. For the part where you got -6, all you have to do is multiply the constants together. -1 * 3 = -3

OpenStudy (anonymous):

wait why -1? why not 1*-3?

OpenStudy (anonymous):

even though its the same. The 3 is - not the 1.